OpenStudy (anonymous):

Write each of the following numerals in base 10. For base twelve, T and E represent the face values ten and eleven, respectively. a. 111 five b. 11101 two c. 31T twelve

OpenStudy (anonymous):

a. 1*5^2 + 1*5^1 + 1*5^0 = 25 + 5 + 1 = 31 b. 1*2^4 + 1*2^3 + 1*2^2 + 0*2^1 + 1 = 16 + 8 + 4 + 1= 29 c. 3*12^2 + 1*12 + 10 = 432 + 12 + 10 = 454